Given that K = 3.83 for the reaction C(s,graphite)+ CO2(g) ↔ 2 CO(g) at 606 °C, determine the Gibbs reaction energy (in kJ mol-1) at 606 °C.
Answer #1

et coz 200 K=3.83 T= 606°C = 813 +606 = 879 K on 2 - RT enk; R =8-314 molk 2 – 8,3148879x en 3083 = -9814 66= -9.81 ko ay Ga=